OpenStudy (anonymous):

x=1670 ft the geometric mean of two numbers is 8. one of the numbers is 2. find the other numbers **********metal*************

OpenStudy (anonymous):

\[\sqrt{2x}=8\] solve for \(x\)

OpenStudy (anonymous):

can u solve for x please @satellite73

Directrix (directrix):

If 8 is the geometric mean between two numbers, 2 and x, then the following equation is true. 2 is to 8 as 8 is to x. 2/8 = 8/x Cross Multiply 2x = 64 x = ? @mary_arevalo_boo